ok here is my question. help would be greatly appreciated!!

Rearrange 4y - x = 8 so that it is the form y = mx = b

Answer by bucky(2189)

Given: 4y - x = 8
.
Rearrange it so it is in the form y = mx + b
.
Notice that the form you are to get it in has only a "y" on the left side. The "x" term and
the constant "b" are on the right side.
.
Now look at the given equation. The left side has a -x. It must be removed from the left
side. You can do this by adding a +x to both sides of the equation. On the left side the
-x and the +x that is being added cancel each other out. This and the addition of +x to the right
side makes the equation become:
.
4y = +x + 8
.
Since you are to have only "y" on the left side, you need to divide all terms on both sides
by 4 which is the multiplier of x. This division of both sides results in:
.
4y/4 = (1/4)*x + 8/4
.
and this simplifies to:
.
y = (1/4)*x + 2
.
This is the form you were to get it in. Notice that m = (1/4) and b = + 2
.
Hope this helps you to understand how to rearrange equations to get them into a form that you
need.